Session: Jun 2007 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Wed Jun 20 18:37:16 2007) Success :
S-band RFI. A few bad scan_checks. Otherwise no known problems.

Effelsberg (A. Polatidis at Thu Jun 21 08:19:38 2007) Minor failures :
Lost Day 171 0130-0318, 0400-0541 due to thunderstorms, day 170 1004-1035 due to problem with a diskpack. Day 171, 1436, problems with the subreflector mount, fixed after resetting within a couple of minutes.

Effelsberg (A. Polatidis at Thu Jun 21 12:59:40 2007) Minor failures :
An additional note: The fmout-gps fmout-gps reading jumped from 11.1E-06 sec to 7.9 E-06 sec, between the end of experiment F07SX1 and the start of EC025A.

Medicina (Andrea Orlati at Thu Jun 28 08:11:36 2007) Success :
Nothing relevant to point out.